I dont understand this problem,how do I go about solving 2/5k=22?
Answer by nyc_function(2741)

2/5k = 22
Multiply both sides by 5k.
After doing so, we get 2 = 110k.
We now divide both sides by 110 to find k.
2/110 = k
1/55 = k